# Spirals, rings, and vortices shaped by shadows in protoplanetary disks: from radiative hydrodynamical simulations to observable signatures

### v1

@@ Abstract
Numerous protoplanetary disks exhibit shadows in scattered light observations. These shadows are typically cast by misaligned inner disks and are associated with observable structures in the outer disk such as bright arcs and spirals. Investigating the dynamics of the shadowed outer disk is therefore essential in understanding the formation and evolution of these structures. We carry out twodimensional radiation hydrodynamics simulations that include radiative diffusion and dust–gas dynamics to study the formation of substructure in shadowed disks. We find that spiral arms are launched at the edge of each shadow, permeating the entire disk. The local dissipation of these spirals results in an angular momentum flux, opening multiple gaps and leading to a series of concentric, regularly-spaced rings We find that ring formation is favored in weakly turbulent disks where dust growth is taking place. These conditions are met for typical class-II disks, in which bright rings should form well within a fraction of their lifetime ($\sim$0.1–0.2 Myr). For hotter disks gap opening is more efficient, such that the gap edges quickly collapse into vortices that can appear as bright arcs in continuum emission before decaying into rings or merging into massive, long-lived structures. Synthetic observations show that these structures should be observable in scattered light and millimeter continuum emission, providing a new way to probe the presence of substructure in protoplanetary disks. Our results suggest that the formation of rings and gaps is a common process in shadowed disks, and can explain the rich radial substructure observed in several protoplanetary disks.
@@ 1 Introduction
At the same time, such vigorous spiral activity would provide an efficient mechanism to redistribute angular momentum within the disk. The nonlinear damping of spiral arms can drive a local angular momentum flux into the disk due to them steepening into shocks (e.g., Rafikov, 2002) at the interface between the shadowed and illuminated regions, or due to radiative cooling (Miranda & Rafikov, 2020; Zhang & Zhu, 2020; Ziampras et al., 2023). This process essentially acts to remove material from the location where damping occurred, ultimately leading to the formation of a gap if the angular momentum flux is both sustained long enough at a given radial location and strong enough to overcome the gap-refilling effects of turbulence and local pressure gradients (Crida et al., 2006). Due to the long timescales involved in the gap-opening process, however, the possibility of gap opening due to the spiral arms excited in shadowed disks has not been explored in detail (but see Su & Bai, 2024, whose work was published during the preparation of this manuscript).
@@ 2.3 Shadows due to a misaligned inner disk
Similar to Montesinos et al. (2016), we consider a small, misaligned inner disk casting a shadow on the “main” disk that we actually evolve in our simulations. We assume that the inner disk extends between $R\in[0.2,4]$ au and is tilted by $30^{\circ}$ with respect to the $xy$ plane about the $y$ axis, while the main disk extends beyond 5 au. The two disks otherwise share the same physical properties (see Sec. 2.2), and we apply an exponential taper in the radial direction for both disks for a smooth transition between the two. A detailed description of our two-disk model is provided in Appendix A.
@@
Figure 1: Schematic of the disk configuration. The inner disk is tilted by $30^{\circ}$ with respect to the $xy$ plane and casts a shadow on the main disk.
@@
Here, $\Delta r$ is the radial cell width, and $\rho_{\text{d}}^{\text{small}}$ is the volume density in small dust, assumed to be perfectly coupled to the gas and therefore given by $\rho_{\text{d}}^{\text{small}}(r,\theta)=\varepsilon(1-X)\rho_{\text{g}}$, where $\rho_{\text{g}}$ is the gas density. We set $X=0.9$ and $\varepsilon=0.01$.
@@ 2.4 Numerical setup
Our grid extends radially with a logarithmic spacing between $R\in[6,75]\,\text{au}=[0.2,2.5]\,R_{0}$ with a reference radius of $R_{0}=30$ au and covers the full azimuthal extent $\phi\in[0,2\pi]$ with $N_{r}\times N_{\phi}=576\times 1536$ cells. This grid resolution translates to approximately 12 cells per gas scale height in either direction at $R=R_{0}$ for $h=0.05$. Our disk is initialized according to Eq. (8), with $u_{R}=0$ for both gas and dust and
@@ 3.1 Disk structure
This behavior supports our hypothesis that the constant generation of spiral waves due to the shadows drives an AMF into the disk once the waves dissipate, ultimately leading to the formation of rings. We investigate further by computing the AMF in the disk, defined as (e.g., Miranda & Rafikov, 2020)
@@ 5.1 RADMC-3D setup
It is important to note here that the temperatures computed by RADMC-3D are equilibrium temperatures, in contrast to those used in the dynamic calculations. As a result, RADMC-3D cannot capture the asymmetric azimuthal variation of temperature found in the hydrodynamical models around the shadow. We highlight this difference in Appendix C.
@@ 5.2 Synthetic images
Figure 16: Synthetic images for our fiducial model with $h_{0}=0.05$ (top two rows) as well as $h_{0}=0.07$ and $h_{0}=0.1$ (third and fourth row) at 1.3 mm (leftmost column) and 1.65 $\mu$m (remaining panels) for different disk inclinations and azimuthal rotations. The top row zooms into the (unresolved) inner region at $R\lesssim 18$ au, marked by a box in each panel of the second row. Rings and shadows are visible throughout the entire disk for all values of $h_{0}$. For better visibility on the same scale, the fluxes are divided by a factor of 10, 3, and 10 in the first, third, and fourth row, respectively. A solid white circle marks the beam size of 40 mas, and a dashed circle the coronagraph of radius 100 mas. We use an arcsinh stretch for the color scale of the 1.3 mm maps.
@@
In the second row of Fig. 16, which shows the full disk structure, the dust continuum emission features both a series of bright rings as well as a slightly darker lane corresponding to the shadow cast by the misaligned inner disk. While both the rings and the shadow are also visible in the scattered-light images, the shadowed wedge is more pronounced while the rings are less so. The top row of the same figure shows zoomed-in views of the inner disk regions, highlighting the misaligned inner disk and the bright, directly illuminated inner rim of the outer disk ($R\lesssim 18$ au). However, the structure of this inner region is blurred due to its angular size being comparable to the beam size.
In the third and fourth row of Fig. 16 we show the synthetic images for models with $h_{0}=0.07$ and $h_{0}=0.1$, respectively, similar to the second row of the same figure. As expected, the vortices that form in both models are clearly visible as bright arcs in both mm emission and scattered light due to their dust-trapping properties.
Furthermore, a weak, wide shadow can be observed behind the massive vortex for $h_{0}=0.1$. This happens due to the vortex acting as an optically thick obstacle, casting a shadow on the outer disk. At the same time, however, the dark lane due to the shadow cast by the misaligned inner disk is not visible for $h_{0}=0.1$ and is only faintly visible for $h_{0}=0.07$. We speculate that this happens because the inner disk is spread out over a larger solid angle for larger $h_{0}$ and therefore less optically thick in the radial direction, resulting in a smoother transition between the shadowed and illuminated regions that is further blurred by radiative diffusion in the azimuthal direction.
@@
Figure 17 shows the synthetic ALMA image for our fiducial model with $h_{0}=0.05$, with the left panel corresponding to $t=1000\,P_{0}$ as in all panels shown so far in this section. While the rings are quite faint, they are still visible to the naked eye without requiring any post-processing. The shadow cast by the misaligned inner disk is visible as well. To highlight the time-sensitive nature of the gap opening process, we also show the synthetic ALMA image at $t=2000\,P_{0}$ in the right panel of the same figure, where the rings are now clearly visible as gap opening progresses and more dust accumulates in the resulting pressure bumps.
For completion, we also show the synthetic ALMA images for the models with $h_{0}=0.07$ and $h_{0}=0.1$ in Fig. 18. In these cases, vortices are clearly visible for both models, but the arcs of dust corotating with the vortices for $h_{0}=0.07$ as well as the rings interior to the vortex for $h_{0}=0.1$ are no longer visible by eye.
Figure 17: Synthetic ALMA image at 1.3 mm for our fiducial model, configured to mimic an observation of the system HD 163296 using SIMIO-continuum. Left: $t=1000\,P_{0}\approx 0.16$ Myr. While faint, rings and shadows are both visible by eye. Right: $t=2000\,P_{0}\approx 0.32$  Myr. The rings are now clearly visible as gap opening progresses.
Figure 18: Same as Fig. 17 for models with $h_{0}=0.07$ (left) and $h_{0}=0.1$ (right) at $t=1000\,P_{0}$. Vortices are clearly visible, but fainter features such as rings and arcs corotating with the vortices can no longer be seen by eye.
@@ 6.2 Effect of a moving shadow
In the case of a precessing shadow, Montesinos & Cuello (2018) showed that the emerging spiral structure resembles planet-like signatures, with the radial location of the spiral-launching point being a function of the precession direction and frequency of the shadow. A radial profile of the cooling timescale complicates the picture, as forming radial substructure would require an appropriate balance between the cooling, shadow precession, and orbital/shadow crossing timescales (see, e.g., Su & Bai, 2024).
6.3 Relating our findings to Su & Bai (2024)
@@ Acknowledgments
AZ would like to thank Xuening Bai, Ondřej Chrenko, Dhruv Muley, Giovanni Rosotti, and Prakruti Sudarshan for helpful suggestions and fruitful conversations. This research utilized Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT (http://doi.org/10.5281/zenodo.438045). This work was performed using the DiRAC Data Intensive service at Leicester, operated by the University of Leicester IT Services, which forms part of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K000373/1 and ST/R002363/1 and STFC DiRAC Operations grant ST/R001014/1. DiRAC is part of the National e-Infrastructure. AZ and RPN are supported by STFC grant ST/T000341/1 and ST/X000931/1. TB acknowledges funding from the European Union under the European Union’s Horizon Europe Research and Innovation Programme 101124282 (EARLYBIRD) and funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 325594231, and Germany’s Excellence Strategy - EXC-2094 - 390783311. MB has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (PROTOPLANETS, grant agreement No. 101002188). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. All plots in this paper were made with the Python library matplotlib (Hunter, 2007). Typesetting was expedited with the use of GitHub Copilot, but without the use of any AI-generated text.
@@ Appendix B Comparison with dynamic radiative transfer
To verify how accurate the time-dependent heating/cooling prescription of this paper is, we carry out a 1D vertical time-dependent radiative transfer calculation, for a column of gas and dust in approximate vertical hydrostatic equilibrium orbiting around the star and periodically passing through a shadow. In this appendix we describe the model, albeit in abbreviated form, and refer to details to an earlier and a near-future paper.

### v2

@@ Abstract
Numerous protoplanetary disks exhibit shadows in scattered light observations. These shadows are typically cast by misaligned inner disks and are associated with observable structures in the outer disk such as bright arcs and spirals. Investigating the dynamics of the shadowed outer disk is therefore essential in understanding the formation and evolution of these structures. We carry out twodimensional radiation hydrodynamics simulations that include radiative diffusion and dust–gas dynamics to study the formation of substructure in shadowed disks. We find that spiral arms are launched at shadow edges, permeating the entire disk. The local dissipation of these spirals results in an angular momentum flux, opening multiple gaps and leading to a series of concentric, regularly-spaced rings. We find that ring formation is favored in weakly turbulent disks where dust growth is taking place. These conditions are met for typical class-II disks, in which bright rings should form well within a fraction of their lifetime ($\sim$0.1–0.2 Myr). For hotter disks gap opening is more efficient, such that the gap edges quickly collapse into vortices that can appear as bright arcs in continuum emission before decaying into rings or merging into massive, long-lived structures. Synthetic observations show that these structures should be observable in scattered light and millimeter continuum emission, providing a new way to probe the presence of substructure in protoplanetary disks. Our results suggest that the formation of rings and gaps is a common process in shadowed disks, and can explain the rich radial substructure observed in several protoplanetary disks.
@@ 1 Introduction
At the same time, such vigorous spiral activity would provide an efficient mechanism to redistribute angular momentum within the disk. The nonlinear damping of spiral arms can drive a local angular momentum flux into the disk due to them steepening into shocks (e.g., Rafikov, 2002) at the interface between the shadowed and illuminated regions, or due to radiative cooling (Miranda & Rafikov, 2020a; Zhang & Zhu, 2020; Ziampras et al., 2023). This process essentially acts to remove material from the location where damping occurred, ultimately leading to the formation of a gap if the angular momentum flux is both sustained long enough at a given radial location and strong enough to overcome the gap-refilling effects of turbulence and local pressure gradients (Crida et al., 2006). Due to the long timescales involved in the gap-opening process, however, the possibility of gap opening due to the spiral arms excited in shadowed disks has not been explored in detail (but see Su & Bai, 2024, whose work was published during the preparation of this manuscript).
@@ 2.3 Shadows due to a misaligned inner disk
Similar to Montesinos et al. (2016) and cuello-etal-2019, we consider a small, misaligned inner disk casting a shadow on the “main” disk that we actually evolve in our simulations. We assume that the inner disk extends between $R\in[0.2,4]$ au and is tilted by $30^{\circ}$ with respect to the $xy$ plane about the $x$ axis, while the main disk extends beyond 5 au. The two disks otherwise share the same physical properties (see Sec. 2.2), and we apply an exponential taper in the radial direction for both disks for a smooth transition between the two. A detailed description of our two-disk model is provided in Appendix A.
@@
Figure 1: Schematic of the disk configuration with slices along the $x$ (top) and $y$ axes (bottom). The inner disk is tilted about the $x$ axis by $30^{\circ}$ with respect to the $xy$ plane and casts a shadow on the main disk.
@@
Here, $\Delta r$ is the radial cell width, and $\rho_{\text{d}}^{\text{small}}$ is the volume density in small dust, assumed to be perfectly coupled to the gas and therefore given by $\rho_{\text{d}}^{\text{small}}(r,\theta)=\varepsilon(1-X)\rho_{\text{g}}$, where $\rho_{\text{g}}$ is the gas density (see Eq. (22)). We set $X=0.9$ and $\varepsilon=0.01$.
@@ 2.4 Numerical setup
Our grid extends radially with a logarithmic spacing between $R\in[6,75]\,\text{au}=[0.2,2.5]\,R_{0}$ with a reference radius of $R_{0}=30$ au and covers the full azimuthal extent $\phi\in[0,2\pi]$ with $N_{r}\times N_{\phi}=576\times 1536$ cells. This grid resolution translates to approximately 12 cells per gas scale height in either direction at $R=R_{0}$ for $h_{0}=0.05$. Our disk is initialized according to Eq. (8), with $u_{R}=0$ for both gas and dust and
@@ 3.1 Disk structure
This behavior supports our hypothesis that the constant generation of spiral waves due to the shadows drives an AMF into the disk once the waves dissipate, ultimately leading to the formation of rings. We investigate further by computing the AMF in the disk, defined as (e.g., Miranda & Rafikov, 2020a)
@@ 5.1 RADMC-3D setup
It is important to note here that the temperatures computed by RADMC-3D are equilibrium temperatures, in contrast to those used in the dynamic calculations. As a result, RADMC-3D cannot capture the asymmetric azimuthal variation of temperature found in the hydrodynamical models around the shadow. We highlight this difference in Appendix C. Given that the azimuthally averaged midplane temperature profiles obtained from RADMC-3D are remarkably similar to those from the hydrodynamical models, we replace the RADMC-3D output of the temperature structure within ±$\pm H/2$ about the midplane with the hydrodynamical temperature structure. In doing so, our models can also capture the azimuthal variations in temperature (see e.g., Fig. 5) when computing synthetic images in mm emission. We believe that applying this correction to the disk surface would not necessarily produce more accurate results, as our models are 2D, but instead refer the reader to Zhang & Zhu (2024), who computed scattered light images for their (3D) models.
@@ 5.2 Synthetic images
Figure 16: Synthetic images for our fiducial model with $h_{0}=0.05$ (top two rows) as well as $h_{0}=0.07$ and $h_{0}=0.1$ (third and fourth row) at 1.3 mm (leftmost column) and 1.65 $\mu$m (remaining panels) for different disk inclinations and azimuthal rotations. The top row zooms into the (unresolved) inner region at $R\lesssim 18$ au, marked by a box in each panel of the second row. Rings and shadows are visible throughout the entire disk for all values of $h_{0}$. For better visibility on the same scale, the fluxes are rescaled by a factor shown on the bottom right corner of each panel. A solid white circle marks the beam size of 40 mas, and a dashed white circle the coronagraph of radius 100 mas. A dashed green circle marks the outer edge of our hydrodynamical domain.
@@
In the second row of Fig. 16, which shows the full disk structure, the dust continuum emission features both a series of bright rings as well as a faint azimuthal variation in brightness due to the shadow cast by the misaligned inner disk. In the remaining panels of the same row, corresponding to the scattered-light images, the shadowed wedge is clearly visible but the rings are not. The top row of the same figure shows zoomed-in views of the inner disk regions, highlighting the directly illuminated inner rim of the outer disk ($R\lesssim 18$ au). However, the structure of this inner region is blurred due to its angular size being comparable to the beam size.
In the third and fourth row of Fig. 16 we show the synthetic images for models with $h_{0}=0.07$ and $h_{0}=0.1$, respectively, similar to the second row of the same figure. As expected, the vortices that form in both models are clearly visible as bright arcs in mm emission, and can still be faintly observed in scattered light, but the rings are too faint to be observable. This is a result of the very weak contrast in $\mu$m-sized grains, which largely trace the gas and therefore show $\Delta\Sigma/\Sigma\lesssim 5\%$. The shadows grow progressively wider with increasing $h_{0}$ due to the larger scale height of the inner disk, and a dark lane can even be observed in mm emission due to the stark temperature contrast between the shadowed and illuminated regions.
@@
Figure 17 shows the synthetic ALMA image for our fiducial model with $h_{0}=0.05$, with the left panel corresponding to $t=1000\,P_{0}$ as in all panels shown so far in this section. While the contrast is rather weak, rings and gaps are visible to the naked eye without requiring any post-processing. The shadow cast by the misaligned inner disk is, however, not discernible by eye. To highlight the time-sensitive nature of the gap opening process, we also show the synthetic ALMA image at $t=2000\,P_{0}$ in the right panel of the same figure, where the rings are now clearly visible as gap opening progresses and more dust accumulates in the resulting pressure bumps.
For completion, we also show the synthetic ALMA images for the models with $h_{0}=0.07$ and $h_{0}=0.1$ in Fig. 18. In these cases, vortices are clearly visible for both models, but the arcs of dust corotating with the vortices for $h_{0}=0.07$ as well as the rings interior to the vortex for $h_{0}=0.1$ are no longer visible by eye. For $h_{0}=0.1$, a dark lane due to the shadow cast by the misaligned inner disk is clearly visible in the synthetic ALMA image, as the shadow is wider and more pronounced for larger aspect ratios.
Figure 17: Synthetic ALMA image at 1.3 mm for our fiducial model (see second row of Fig. 16), configured to mimic an observation of the system HD 163296 using SIMIO-continuum. Left: $t=1000\,P_{0}\approx 0.16$ Myr. While faint, rings and gaps are visible by eye. Right: $t=2000\,P_{0}\approx 0.32$  Myr. The rings are now clearly visible as gap opening progresses. Colors are rescaled with an arcsinh stretch.
Figure 18: Same as Fig. 17 for models with $h_{0}=0.07$ (left) and $h_{0}=0.1$ (right) at $t=1000\,P_{0}$. Vortices are clearly visible, but fainter features such as rings and arcs corotating with the vortices are less discernible by eye.
@@ 6.2 Effect of a moving shadow
In the case of a precessing shadow, Montesinos & Cuello (2018) showed that the emerging spiral structure resembles planet-like signatures, with the radial location of the spiral-launching point being a function of the precession direction and frequency of the shadow. A radial profile of the cooling timescale complicates the picture, as forming radial substructure would require an appropriate balance between the cooling, shadow precession, and orbital/shadow crossing timescales.
6.3 Comparison to related studies
cuello-etal-2019 were first to investigate the trapping of dust particles in spirals driven by shadows in a framework similar to ours, but their models did not feature rings or gaps. Instead, the spiral structures presented in their work are much brighter in 1.3 mm emission, contrary to our models where the spirals are quickly overshadowed by the forming rings or vortices. All of the above can be attributed to several different reasons.
For one, the omission of in-plane radiative diffusion in cuello-etal-2019 renders the disk less radiatively efficient, as the in-plane cooling channel (which can be much more efficient in the presence of temperature gradients) is unavailable (Miranda & Rafikov, 2020b; Ziampras et al., 2023). This results in a longer cooling timescale and therefore slower development of the gap opening process. In combination with their significantly shorter integration timescale of $10^{4}$ yr—10 orbits at 100 au, their reference radius—it is not surprising that the only prominent feature in their models is the spiral structure, which develops very early on.
Furthermore, the brightness of the spirals in mm emission, which primarily depends on the local dust-to-gas ratio, heavily depends on the background dust model. cuello-etal-2019 used a dust size distribution with $a_{\text{d}}^{\text{max}}=10$ cm, or $\mathrm{St}^{\text{max}}\sim 2$ at 100 au, such that a substantial fraction of dust grains lies in the regime $\mathrm{St}\approx 0.1$–1 and can very efficiently accumulate around the spiral arms. The fact that the dust particles were treated using post-processing can further exaggerate this effect, as dust–gas dynamics are missed. In contrast, our models feature mm grains with $a_{\text{d}}^{\text{big}}=1$ mm or $\mathrm{St}\approx 0.01$ at $R_{0}=30$ au, which do not experience such efficient accumulation around spirals but will nevertheless be trapped in the resulting pressure bumps.
Nevertheless, the studies by both Montesinos et al. (2016) and cuello-etal-2019 laid the groundwork in probing the observational signatures of spirals induced in shadowed disks, and in particular our results relate to cuello-etal-2019, where the central focus was the dynamics of the dust component. Our results, while seemingly qualitatively different, are in fact fully consistent with theirs once the differences in parameter space and model complexity are taken into consideration.
@@ 6.3 Comparison to related studies
In addition, recently Zhang & Zhu (2024) showed with threedimensional radiative hydrodynamics simulations that the spiral-driving mechanism due to a shadow is robust and can excite substantial angular momentum transport as well as bright features in scattered-light observations. While their work focused on the spiral formation process and its observable signatures, it is worth pointing out that their models also featured gaps and vortices (see Fig. 2 therein) due to the same mechanism studied in this work. Combined with the recent linear theory calculations by zhu-etal-2024, it is clear that shadow-induced spirals can very efficiently transport angular momentum and drive substructure in protoplanetary disks.
All in all, our findings are in very good agreement with previous and current studies, with a key difference being the inclusion of dust hydrodynamics. In this context, the latter enables more consistent dust–gas interaction both via momentum exchange and the determination of the opacity, as well as synthetic observations of mm emission with ALMA. Further work is needed to understand dust–gas dynamics in a threedimensional model, given that the excitation of spirals can drive substantial vertical motion (Zhang & Zhu, 2024) which can set the dust scale height and therefore overall cooling efficiency.
@@ Acknowledgments
The authors thank the anonymous referee for a careful and constructive report that helped identify errors related to the RADMC-3D models and greatly improved the quality of the manuscript. AZ would like to thank Xuening Bai, Ondřej Chrenko, Dhruv Muley, Giovanni Rosotti, and Prakruti Sudarshan for helpful suggestions and fruitful conversations. This research utilized Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT (http://doi.org/10.5281/zenodo.438045). This work was performed using the DiRAC Data Intensive service at Leicester, operated by the University of Leicester IT Services, which forms part of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K000373/1 and ST/R002363/1 and STFC DiRAC Operations grant ST/R001014/1. DiRAC is part of the National e-Infrastructure. AZ and RPN are supported by STFC grant ST/T000341/1 and ST/X000931/1. TB acknowledges funding from the European Union under the European Union’s Horizon Europe Research and Innovation Programme 101124282 (EARLYBIRD) and funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 325594231, and Germany’s Excellence Strategy - EXC-2094 - 390783311. MB has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (PROTOPLANETS, grant agreement No. 101002188). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. All plots in this paper were made with the Python library matplotlib (Hunter, 2007). Typesetting was expedited with the use of GitHub Copilot, but without the use of any AI-generated text.
@@ Appendix B Comparison with dynamic radiative transfer
To verify how accurate the time-dependent heating/cooling prescription of this paper is, we carry out a 1D vertical time-dependent radiative transfer calculation, for a column of gas and dust in approximate vertical hydrostatic equilibrium orbiting around the star and periodically passing through a shadow. In this appendix we describe the model, albeit in abbreviated form, and refer to Dullemond et al. (2002) and Dullemond et al. (2022) for details.